OpenStudy (anonymous):

Find the derivative of y=(5x+10)^5/(x-3)^2

OpenStudy (kenljw):

y=(5x+10)^5/(x-3)^2 N'(x) = 5 (5x+10)^4 5 = 25 (5x+10)^4 chain rule D'(x) = 2 (x-3) y' = (N'(x) D(x) -N(x) D'(x))/D(x)^2 quotient rule

OpenStudy (anonymous):

thank you! :)